OpenStudy (anonymous):

At Cold Marble Ice Cream, you create your own ice cream flavor by choosing from a list of 10 "add-ins" that can be mixed into vanilla ice-cream. If you can pick one add-in, any combination of two add-ins, or any combination of three add-ins, how many different flavors are possible?

OpenStudy (anonymous):

\[10+10\times 9+10\times 9\times 8\]

OpenStudy (anonymous):

you got that?

OpenStudy (anonymous):

I don't get any of the answers when I do this or the other one..